OpenStudy (anonymous):

can anyone help? sketch the line 3x + 4y = 24, indicating the intercepts on both axes

OpenStudy (anonymous):

solve the equation for y: 4y=-3x+24 y=-3/4x+6 The line crosses the y axis at 6 (y-intercept is 0,6). Another way to get this is y=-3/4(0)+6 y= 6 (0,6) To get the x intercept, solve the equation when y=0. 0=-3/4x+6 -6=-3/4x -8 = x x intercept is (-8,0)

OpenStudy (anonymous):

thanks so much!